### v1

# OptiMUS-0.3

[…]
OptiMUS-0.3
[…]
OptiMUS-0.3: Using large language models to model and solve optimization problems at scale
[…]
Ali AhmadiTeshnizia , Wenzhi Gaob, Herman Brunborgb Shayan Talaeia, Madeleine Udella,b
[…]
Optimization problems are pervasive in sectors from manufacturing and distribution to healthcare. However, most such problems are still solved heuristically by hand rather than optimally by state-of-the-art solvers because the expertise required to formulate and solve these problems limits the widespread adoption of optimization tools and techniques. We introduce a Large Language Model (LLM)-based system designed to formulate and solve (mixed integer) linear programming problems from their natural language descriptions. Our system is capable of developing mathematical models, writing and debugging solver code, evaluating the generated solutions, and improving efficiency and correctness of its model and code based on these evaluations. OptiMUS-0.3 utilizes a modular structure to process problems, allowing it to handle problems with long descriptions and complex data without long prompts. Experiments demonstrate that OptiMUS-0.3 outperforms existing state-of-the-art methods on easy datasets by more than 12% and on hard datasets (including a new dataset, NLP4LP, released with this paper that features long and complex problems) by more than 8%.
[…]
Large Language Models, Optimization Modeling, AI Agents, Mixed Linear Integer Programming, Optimization Tools, Human-Computer Interaction
1 Introduction
Optimization problems are common in domains from operations and economics to engineering and computer science. Major advances in optimization algorithms over the last several decades have led to reliable and efficient optimization methods for a wide variety of structured optimization problems, including linear programming (LP) and mixed-integer linear programming (MILP), among many others. Unfortunately, optimization modeling — transforming a domain problem into a mathematical optimization problem — still requires expert knowledge. According to a recent survey, 81% of Gurobi’s commercial solver users have advanced degrees, with 49% of them holding a degree in operations research (Gurobi Optimization 2023). This expertise gap prevents many organizations from using optimization, even when it could significantly improve their operations. Examples include inventory management in supermarkets, patient operations in hospitals, transportation policies in small municipalities, energy management in local solar farms, and operations in small businesses or NGOs (Saghafian et al. 2015, Aastrup and Kotzab 2010, Yao et al. 2020, Shakoor et al. 2016). Automating optimization modeling would allow enterprises to improve efficiency using optimization techniques, even when they cannot afford to access optimization experts.
Large language models (LLMs) offer a promising way to automate optimization modeling, and increase the reach of the powerful algorithms developed by the operations research community. However, these LLM models are still in their infancy and suffer from major flaws that prevent deployment in important applications. Most problematically, it is difficult to ensure correctness and completeness of the LLM output. We highlight four challenges:
  • Long problem descriptions. Realistic optimization problems can be exceedingly long and complex: for example, the documentation for the energy system problem described in Liu et al. (2023b) is 60 pages long. Unfortunately, LLMs have a limited context size, and even long-context models perform less well as the input context grows Liu et al. (2023b), Levy et al. (2024). Consequently, LLMs tend to make more mistakes as the length of the problem description increases and perform poorly on complex problems. How can we design an LLM-based optimization modeling system that keeps track of each individual constraint and term in the objective and ensures the model is complete and correct?
  • Large problem data. The specification of an optimization problem often involves large amounts of data, such as customer attributes or sales of goods. Efficient formulations that ensure fast solve times are essential to handle large problem data and competing with expert-level solutions. Previous approaches to optimization modeling using LLMs, which pass numerical data to the LLM directly and use simple formulations, are thus restricted to the simplest of toy problems.
  • Hallucination. LLMs are known to hallucinate: they may produce answers that sound reasonable, but are incorrect. In the context of optimization, the generated solver code may hallucinate constraints that incorrectly model the problem, or it may hallucinate API calls that do not exist, resulting in code that cannot run. It is especially challenging to verify whether the solution is correct, supposing the code runs without error. For instance, if the solver claims the solution is unbounded, perhaps a constraint has been accidentally omitted from the formulation.
  • Bad models. The solve time for an optimization problem can depend on the particular modeling formulation chosen, and on how the structure of the problem is communicated to the solver. Optimization experts spend much of their effort modeling the problem to enhance the efficiency of the solution method. A challenge for LLMs is to produce not just an accurate model, but good code that solves the problem quickly.
This paper studies the feasibility of using LLMs to expand the reach of operations research techniques and analytic methods for decisionmaking by studying the performance of OptiMUS, an automated system we have developed that uses LLMs combined with traditional solvers to model and solve (mixed integer) linear programs (MILPs). OptiMUS advances the agenda of operations research when viewed as a meta-analytic tool to improve decisionmaking. We have chosen to begin with MILP modeling as MILP is one of the more classic, important, and widely used tools in the operations research toolbox. Moreover, examples of MILP problems, models, and solutions are plentiful and readily available. Our hope is that by making powerful solvers easily accessible to domain experts, OptiMUS can empower better, faster decisions that reduce costs and improve outcomes. Our studies of OptiMUS demonstrate both the promise and pitfalls of the LLM approach to quantitative decisionmaking. More broadly, through ablation studies, this paper clarifies the information flow needed to perform optimization modeling: in other words, it provides empirical evidence for what kind of education is effective to teach (LLMs) optimization modeling.
We call the version of OptiMUS studied in this paper OptiMUS-0.3 to distinguish it from previous versions that have been described in the arXiv draft AhmadiTeshnizi et al. (2023) and the conference publication Ahmaditeshnizi et al. (2024). Compared to these previous versions, OptiMUS-0.3 uses an LLM-based pipeline instead of the more flexible but error-prone agent-based system of OptiMUS-0.2. Our findings indicate that the increased flexibility in OptiMUS-0.2 led to more errors, outweighing the benefits at the current LLM capabilities. Additionally, nondeterministic errors made maintaining and improving the system challenging. Unlike OptiMUS-0.1, OptiMUS-0.3 uses a modular approach, using only relevant information for each constraint. OptiMUS-0.3 also incorporates Retrieval-Augmented Generation (RAG) and error correction, which were absent in both previous versions.
Our contributions in this paper are as follows:
  • We identify 18 important (non-exclusive) types of MILP problems, including scheduling, routing, production planning, and inventory management, which we use as labels for optimization problems. We curate a new dataset, NLP4LP, a comprehensive open-source dataset of 252 optimization problems that ensures adequate coverage of all these problem types, by combining problems from existing datasets with new problems. This dataset also includes problems that range in difficulty and description length, including real-world problems that are an order of magnitude longer that the problem instances in other MILP modeling datasets. Table 1 compares NLP4LP to existing datasets and Section 3 describes NLP4LP.
  • We develop a modular, LLM-based agent to model and solve optimization problems, which we call OptiMUS-0.3. OptiMUS-0.3 employs a connection graph that allows it to process each constraint and objective independently, which allows the system to solve problems with long descriptions and large data files without excessively long prompts.
  • We develop several new modules designed to improve the performance of OptiMUS-0.3 and study their impact, including:
    1. 1.
      Self-reflective error correction. We ask OptiMUS to assess its own confidence in its output, and fall back to a more powerful LLM or to user feedback when OptiMUS is unsure its answer is correct.
    2. 2.
      Retrieval-augmented generation. We search through a database of constraints and their formulations to find ones similar to the task at hand. We include a few of these as examples in the prompt which OptiMUS passes to the LLM.
    3. 3.
      Advanced optimization modeling techniques. We teach OptiMUS advanced optimization modeling techniques by prompting it to identify important structures in the problem, including structure of problems and solver features such as special ordered sets, and to use those structures to model and solve the problem more efficiently. We observe consistent improvement in terms of scalability and solution time after proper structure is identified.
  • We perform ablation studies of the components of our framework to demonstrate which elements improve performance and which can actually degrade performance. We find OptiMUS-0.3 beats the previous state-of-the-art methods on existing datasets by over 12% and on our more challenging test dataset by 8%.
2 Related Work
Optimization problems are mathematically defined by an objective function and a set of constraints. For example, an MILP can be written mathematically as
$\underset{\{x_{j}\}}{\text{minimize}}\quad$ $\sum_{j=1}^{n}c_{j}x_{j}$
subject to $\sum_{j=1}^{n}a_{ij}x_{j}{~{}(\leq,=,\geq)~{}b_{i}},i=1,\ldots m$
$l_{j}\leq x_{j}\leq u_{j},j=1,\ldots,n$
$x_{j}\in\mathbb{Z},j\in\mathcal{I}$
An optimization workflow consists of 1) formulating an optimization problem in mathematical form by identifying its objective and constraints, and then 2) solving the realization of problem from real data, generally using code that calls an optimization solver.
Formulation is often a challenging task even for experts in optimization. Different formulations can lead to significantly different solving times and enable the use of different solvers or solution techniques (Boyd and Vandenberghe 2004). One important skill for an optimization expert is to identify assumptions or relaxations that allow for casting the problem as a well-studied problem type, such as MILP, which enables the use of well-developed optimization solvers. Crafting such an efficient formulation often requires specialized knowledge (Zohrizadeh et al. 2020, Low 2013, Roubíček 2020, Luo et al. 2010, Krarup and Pruzan 1983).
Given the formulation, an optimization expert must choose a solver. Each solver has its own interface and functionalities (Achterberg 2019, Diamond and Boyd 2016, CPLEX User’s Manual 1987). However, the user manuals for these solvers are often hundreds of pages, making them difficult to fully understand and use.
Table 1: A comparison on different aspects of complexity for various datasets. The unit for description length is characters
Dataset Description Length Instances (#MILP) Multi-dimensional Parameters
NL4Opt 518.0 $\pm$ 110.7 1101 (0) $\times$
ComplexOR 497.1 $\pm$ 247.5 37 (12) $\checkmark$
NLP4LP Easy (Ours) 507.2 $\pm$ 102.6 287 (0) $\checkmark$
NLP4LP Hard (Ours) 912.3 $\pm$ 498.2 67 (18) $\checkmark$
Progress in LLMs.
Recent progress in Natural Language Processing (NLP) has led to the development of large language models (LLMs) useful for tasks such as answering questions, summarizing text, translating languages, and coding (OpenAI 2023, Touvron et al. 2023, Chowdhery et al. 2022, Wei et al. 2023, Gao et al. 2023a, Borgeaud et al. 2022). Connections to other software tools extend the reach and accuracy of LLMs, as demonstrated by plug-ins for code writing and execution (Paranjape et al. 2023, Wei et al. 2023). (Yang et al. 2023) use LLMs to directly generate solutions to optimization problems without calling traditional solvers through prompt optimization to improve performance. The approach is limited to small problems since the performance of LLMs degrades as the input context grows, even for explicitly long-context models (Liu et al. 2023a).
Chatbots for Optimization.
In a recent paper, (Chen et al. 2023) developed a chatbot to help users detect and fix infeasible optimization problems expressed in Pyomo code and servers as an AI assistant rather than as a solver. (Li et al. 2023) designed a chatbot to answer natural-language queries about an optimization model. (Alibaba Cloud 2022) also developed a chatbot to facilitate optimization modeling, but there is no public paper or documentation available on it.
Benchmark-driven Optimization Modeling.
More closely related to our approach, Ramamonjison et al (2023) introduced a dataset of 1101 natural language representations of LP problems. They proposed a two-stage mapping from the natural-language representation to the problem formulation using an intermediate representation. Ramamonjison et al (2022) designed a system to simplify and improve the modeling experience for operations research, but did not offer an end-to-end solution. Xiao et al. (2023) presented a multi-agent cooperative framework called Chain of Experts (CoE) to automatically model and program complex operation research (OR) problems, and evaluated it on NL4Opt and another more complex dataset, ComplexOR, introduced in that paper. Concurrent with the preparation of this paper, Tang et al. (2024) introduced a new approach to OR modeling with LLMs that uses fine-tuning on a semi-synthetic dataset rather than prompt optimization or agentic models, which they call ORLM (Operations Research Language Model). The numerical performance of their method is comparable to OptiMUS, but the work is complementary, as the fine-tuning methodology of ORLM could be combined with the agentic framework developed in the present work to improve performance further.
OptiMUS-0.3 is a variation of the previous optimization modeling systems we have built. OptiMUS-0.1 AhmadiTeshnizi et al. (2023) uses a set of unit tests generated by the LLM and revised by the user to evaluate and fix the generated code. If the code does not run or fails the tests it is passed to the LLMs along for the relevant error code for revision until it is fixed. OptiMUS-0.2 Ahmaditeshnizi et al. (2024) uses a structured sequence of LLM agents to effectively model and solve optimization problems. First, the natural language representation of the problem is preprocessed into a structured problem. Then, a team of agents iteratively augments the structured problem with a connection graph, mathematical formulations of each clause, and code for each clause, until the problem is solved. OptiMUS-0.3 uses the modular connection graph from OptiMUS-0.2, the pipeline from OptiMUS-0.1, and new error correction methods to achieve state-of-the-art performance.
Traditional MILP solvers generally benchmark against the MIPLIB benchmark, which offers a diverse collection of MILP problems in standard form. Unfortunately, most of these problems are not associated with a natural-language description, and so cannot be used to study optimization modeling as we do in this paper.
3 Dataset
Resource AllocationProduction PlanningSchedulingBlendingWorkforce ManagementSelectionSizingRoutingInventory ManagementPackagingPricingBalancingPortfolio OptimizationNetwork FlowCoveringCuttingFittingLayout0050505050100100100100150150150150200200200200250250250250300300300300350350350350324324324324179179179179808080805555555537373737363636363333333328282828171717171515151513131313131313131010101088884444333333332222ValuesOperational Areas
Figure 1: NLP4LP covers all common optimization problem types
Manufacturing and ProductionSupply Chain ManagementFood and BeverageTransportation and LogisticsHealthcare and MedicalRetail and E-commerceEnvironmental and SustainabilityAgriculture and ForestryScience and ResearchEnergy and Power SystemsFinance and BankingSports and EntertainmentGovernment and Public SectorEducationHuman ResourcesTelecommunicationsMarketing and MediaAerospace and Defense002020202040404040606060608080808010010010010012012012012014014014014014214214214210210210210278787878727272726868686853535353303030302929292919191919151515151515151515151515131313131212121212121212333333332222ValuesIndustry Sectors
Figure 2: NLP4LP covers most common optimization problem domains
23456780050505050100100100100150150150150ValuesRatings Count
Figure 3: Most problems have around 4 labels.
We introduce NLP4LP, a comprehensive open-source dataset of 344 optimization problems. Our goal in creating NLP4LP is to provide the community with examples they can use to design optimization modeling tools using natural language systems and to assess their quality. It consists of two sub-datasets: 1) An easy dataset which contains only LP problems with short descriptions and scalar parameters, and 2) a hard dataset which contains both LP and MILP problems with longer descriptions and multi-dimensional parametrs (Table 1).
Given the complexity of the task, and the components of each instance (description, sample feasible data, and solution code), gathering high-quality optimization instances is expensive. Unlike traditional ML datasets, NLP4LP does not provide a training set with the same distribution as the test set. As the cost of collecting each instance is high, we have prioritized the task of benchmarking by emphasizing the test set rather than reducing the size of the test set to provide training data. Consequently, NLP4LP is partitioned into 1) a development (dev) set of 22 problems (12 easy and 10 hard) and 2) a test set of 332 problems (277 easy and 55 hard). We designed the partition intending that the dev set be used to develop optimization modeling tools, while the test set be used to evaluate such tools. The dataset is available publicly at https://nlp4lp.vercel.app/
Optimization modeling is used in many domains. To ensure that NLP4LP comprehends a wide variety of use cases, we first create a list of important types of optimization problems, such as scheduling, cutting, routing, blending, and packing, and a list of common application domains such as sports, government, retail, agriculture, and energy. We tag each problem NLP4LP with all relevant labels, and continue to gather instances until the dataset contains at least two instances per label. We collect the instances from textbooks, lecture notes, research papers, and existing datasets Bertsimas and Tsitsiklis (1997b), Williams (2013), Nace (2020), Xiao et al. (2023), Ramamonjison et al (2023) Figure 3 shows that most problems have around four labels, and some have as few as two or as many as eight. Figures 1 and 2 provide histograms showing the number of problem instances for each type and domain; resource allocation problems are the most common.
Real-world problems are often much longer and more complicated than textbook problems. For the test set, we have included ten long real-world problems from scientific papers. These problems are an order of magnitude longer than the instances in the other datasets (in terms of number of variables and constraints).
4 Methodology
This section details the design of OptiMUS-0.3. The workflow of OptiMUS-0.3 is outlined in Algorithm 1. For brevity, the term clause refers to a constraint or objective within the problem.
OptiMUS-0.3 begins by extracting parameters from the problem description (line 2). It then extracts (line 4), and formulates (line 7) the clauses. During clause formulation, the variables, their auxiliary constraints, and a connection graph (a graph that records which variables and parameters appear in each constraint) are also generated. OptiMUS uses this connection graph to ensure consistency of the formulations. The connection graph is key to performance and scalability of OptiMUS, as it allows the LLM to focus only on the relevant context for each prompt, generating more stable results. An example of the constraint formulation step appears in Fig. 4
Figure 4: The formulation process for a single constraint. OptiMUS-0.3 identifies any parameters and variables appearing in the constraint, including new variables that it may need to define. It defines new variables as needed, updates the connection graph which records which constraints use which parameters and which variables, and annotates the constraint with a formulation. (dashed lines represent new connections and variables)
Subsequently, OptiMUS-0.3 generates code snippets for all parameters, clauses, and variables (line 11), assembles these snippets into a single runnable code file (line 12), and iteratively executes and debugs to achieve the desired results (lines 13-15) or until timeout. Error correction (EC) after the extraction and formulation stages catch common errors and improve the solve rate. All prompts used in OptiMUS-0.3 appear in Concluding the reduction.
Algorithm 1 Workflow of OptiMUS-0.3
1:Input: Natural language description of problem $\mathcal{P}$
[…]
3:$\textsc{CorrectErrors}(\textup{params},Q_{params})$
[…]
5:$\textsc{CorrectErrors}(\textup{cons},Q^{(nl)}_{cons})$
6:$\textsc{CorrectErrors}(\textup{obj},Q^{(nl)}_{obj})$
[…]
8:$\textsc{CorrectErrors}(\textup{cons},Q^{(form)}_{cons})$
9:$\textsc{CorrectErrors}(\textup{obj},Q^{(form)}_{obj})$
10:$\textsc{CorrectErrors}(\textup{vars},Q_{vars})$
[…]
13:while Execute(code) throws error e do
[…]
15:end while
16:res $\leftarrow$ Execute(code)
Algorithm 2 CorrectErrors(X, Q)
1:Input: a set of JSON objects $X$, and a list $Q$ of tuples of the form $(q_{llm},q_{user},f)$ where $q_{llm}$ is an LLM prompt template, $q_{user}$ is a user prompt template, and $f$ is a callback function.
2:for $x\in X$ do
3:     for $(q_{llm},q_{user},f)\in Q$ do
4:         $\mathrm{prompt}\leftarrow q_{llm}.\mathrm{format}(x)$
5:         $\mathrm{response},\mathrm{confidence}\leftarrow\mathrm{LLM}(prompt)$
6:         if $\mathrm{confidene}\leq\lambda$ then $\triangleright$ If the LLM is unsure what to do, then ask the user
7:              $\mathrm{user\_prompt}\leftarrow q_{user}.\mathrm{format}(\mathrm{response},x)$
8:              $\mathrm{response}\leftarrow\mathrm{AskUser}(\mathrm{user\_prompt})$
9:         end if
10:         $x\leftarrow f(\mathrm{response},x)$triangleright Update the object
11:     end for
12:end for
13:return $X$
4.1 State
OptiMUS-0.3 manages and modifies the solution using states saved in JSON format. The state consists of the following components:
  • Parameters: OptiMUS-0.3 can choose a symbol for each parameter, infer its shape, and define the parameter if it is not explicitly included in the problem statement. Numerical data from the problem statement is omitted from the parameter list and stored separately for later use, ensuring that the parameter list remains concise and easy to include in future prompts.
  • Clauses: Each clause (objective or constraint) contains a natural language description, a formulation, and a definition code.
  • Variables: Like parameters, each variable has a symbol, a shape, and a definition. Unlike parameters, each variable also has a type, which can be Continuous, Integer, or Binary.
  • Background: The background is a short string that explains the real-world context of the problem. Including this string in every prompt helps improve common sense reasoning.
Figure 5: A completed state for a factory production optimization problem
The state is initialized as an empty object, and is completed step by step throughout the process. An example of a completed state for a factory optimization problem appears in Fig. 5.
4.2 Error Correction
To build a trustworthy and reliable system using LLMs, it is important to mitigate the impact of LLM hallucinations. In the context of optimization modeling, an LLM might confidently generate incorrect parameters, redundant mathematical constraints, or erroneous code. To address this issue, OptiMUS-0.3 uses two main error correction techniques: reflective prompts and confidence-based user feedback.
Figure 6: OptiMUS-0.3 can fix its parameter and variable identification errors when prompted “Is the value of $P$ known or not?”
4.2.1 Reflective prompts
LLMs can often identify and fix their mistakes by using reflective prompts Shinn et al. (2023). We analyzed the most common types of mistakes made by LLMs at each step of the process and designed reflective prompts to address each one. Using these prompts significantly reduces the modeling error rate. The resulting substantial improvement to OptiMUS substantiates the claim that domain-specific knowledge can be used to develop LLM applications that outperform general-purpose LLMs (see Table 3). The reflective prompts used by OptiMUS-0.3 are as follows:
Parameter Extraction.
The model often confuses parameters with variables, misidentifies parameter shapes or misses some parameters. We use the following reflective prompt to correct these errors:
Figure 7: OptiMUS-0.3 can fix parameter identification errors when prompted “Is the value of $P$ known or not?”
  • “Is the value of $P$ known or not? Based on that, is it a parameter or a variable?” (see Fig. 7)
Constraint Extraction.
During constraint extraction, the model sometimes extracts trivial or vague constraints and may regenerate a previously generated constraint in different language. We use the following reflective prompts to correct these errors:
Does constraint $C$ need to be explicitly modeled in the mathematical formulation? (see Fig. 6) Here’s the list of all constraints: $C_{1},C_{2},\ldots,C_{k}$. Are any of them redundant?
Figure 8: OptiMUS-0.3 can fix its constraint modeling errors when prompted “Are units the same for both sides of $C$?
Clause Modeling. Once the clauses are extracted, OptiMUS-0.3 models them, defining variables and auxiliary constraints if needed. Often, the model generates incorrect mathematical formulations for constraints. We use the following reflective prompts to detect and correct these errors:
Are units the same for both sides of $C$? (see Fig. 8) Is $C$ a correct constraint in the context of MILP solvers? (see Fig. 9) Does constraint $C$ involve any variables, or is it only defined over parameters?
Figure 9: OptiMUS-0.3 can fix its constraint modeling errors if by asking Is constraint $C$ a correct constraint in the context of MILP solvers?
Figure 10: OptiMUS-0.3 can fix its parameter and variable identification errors when prompted “Is the value of $P$ known or not?”
OptiMUS-0.3 uses similar reflective prompts to perform error checking for the optimization objective. For variables, we use the following prompt:
Is the value of $V$ known or not? Based on that, is it a parameter or a variable? (see Fig. 10)
4.2.2 Confidence-based user/LLM feedback
In the context of real-world optimization modeling, it is exceeding rare for a natural-language problem description to correspond to an unambiguous MILP formulation. Rather, the process of optimization modeling generally involves many cycles of formulation and user feedback to arrive at a correct model for the problem. OptiMUS-0.3 simulates this process in order to detect errors and improve its solve rate. OptiMUS-0.3 uses a confidence-based reflective method to decide when to request help with its formulation, either from a user (if available) or another (more powerful) LLM; for simplicity, we refer to this partner as the user below.
First, OptiMUS-0.3 scores its confidence in its output:
How confident are you in your response (from 1 to 5)?
If the confidence score is less than 4, OptiMUS-0.3 asks the user to help. The user 11 can decide to either keep the constraint, remove it, or modify it.
Figure 11: Asking GPT-4o to provide feedback when LLama-3-70B’s confidence is low in identifying a constraint.
5 Experiments
In this section, we conduct a comprehensive evaluation of OptiMUS-0.3. We begin by detailing the datasets used in our experiments and showcase the superior performance of OptiMUS-0.3 across these datasets, highlighting its strengths. An ablation study demonstrates the impact of different system components on our results, and a sensitivity analysis probes the internal dynamics of OptiMUS-0.3. We conclude this section by identifying failure cases and potential areas for further improvement.
5.1 Overall Performance
To evaluate the overall performance of OptiMUS, we compare it with standard prompting, Reflexion, and Chain-of-Experts (CoE) Shinn et al. (2023), Xiao et al. (2023). Reflexion is the highest-performing general-purpose framework and CoE is the state-of-the-art method for natural-language optimization modeling. Three main metrics have been used in the literature: accuracy, compilation error (CE) rate, and runtime error (RE) rate. However, a method can generate a totally irrelevant short code that runs, or fix runtime and complication errors by completely removing relevant sections of the code. Hence, we only compare the models’ accuracy. Accuracy is defined as the number of instances correctly solved (An instance is considered as correctly solved only if the code runs successfully and the optimal value is correct. Optimal values are obtained from the dataset or by solving the problems manually). Results are presented in Table 3. OptiMUS-0.3 outperforms all other methods in all datasets by a large margin. This remarkable performance improvement highlights the importance of modularity and structure compared to a single prompt to solve complex problems using LLMs.
5.2 Ablation Study
Table 3 shows the impact of different components, as well as choice of LLM on the performance of OptiMUS. One interesting observation is the performance drop that occurs when less capable LLMs are used instead of GPT-4-o. Optimization tasks need complicated reasoning, and we believe that smaller models are mostly optimized on simpler tasks. Nonetheless, OptiMUS-0.3 can achieve an acceptable performance using the open-source LLaMa3-70B.
We can observe that the constraint extraction error correction step is cruicial to OptiMUS-0.3’s performance. The modeling EC, however, is not as important, mainly because most of the modeling errors can later be fixed in the debugging step (while constraints can only be added in the extraction step).
The LLM feedback step marginally improves the performance of the system. The debugging step, is particularly important for harder problems, given LLMs often make a few small mistakes when they generate code.
Fig. 12 left shows the relation between the performance and the number of debugging steps for different models. In all models, the first few debugging steps increase the performance, but the gain plateaus afterwards. Fig. 12 shows the composition of different failure reasons for easy and hard problems. Clause extraction and clause modeling are the most common reason of failure.
5.3 Failure Cases
To understand its strengths and weaknesses, we analyze the most common reasons why OptiMUS fails. We categorize failure cases into the following groups (Fig. 12):
  • Missing or wrong clauses: OptiMUS generates wrong constraints or objective (e.g., price $\geq$ 0 where price is a parameter), or fails to extract all of the constraints from the description.
  • Incorrect model: OptiMUS tackles the problem with an incorrect mathematical model (e.g., defining binary variables for visiting cities instead of links in TSP).
  • Coding error: OptiMUS does not generate error-free code even after debugging. Coding errors often occur when the LLM is confused by the language used (e.g. a runtime error because of accessing parameter array indices that do not exist).
Understanding and interpreting the problems is also challenging for LLMs, resulting in formulations with missing constraints and wrong constraints. Fine-tuning might improve the performance of LLMs on this task, and is an important direction for future research.
Table 2: OptiMUS-0.3 outperforms the other methods
Easy Hard
Standard 47.3% 33.2%
Reflexion 53.0% 42.6%
CoE 64.2% 49.2%
OptiMUS-0.2 78.8% 68%
OptiMUS-0.3 88.6% 73.7%
Table 3: Ablation studies’ results on OptiMUS-0.3
Easy Hard
Importance of Different Components
w/o Debugging 73.2% 26.7%
w/o Extraction EC 86.7% 60.5%
w/o Modeling EC 83.8% 65.7%
w/o LLM Feedback 88.6% 68.4%
OptiMUS-0.3 88.6% 73.7%
Performance with Different LLMs
LLaMa3-70B 77.3% 53.3%
GPT4-o 88.6% 73.7%
1111222233334444555566660020202020404040406060606080808080100100100100Number of Debugging IterationsPerformance (%)GPT4-o EasyGPT4-o HardLLaMa3-70B EasyLLaMa3-70B Hard
ExtractionModelingCoding0020202020404040406060606080808080100100100100Failure TypeRate(%)EasyHard
Figure 12: Left) More debugging iterations improves the performance. Right) For harder problems, most failures happen during clause extraction. For easier problems, most failures happen during the coding step.
Table 4: Error correction methods can find and fix a large fraction of errors in constraint extraction (left) and constraint modeling (right), while leaving the correct items intact most of the time (a diagonal matrix is good)
Not Modified Modified
Right 219 7
Wrong 9 41
Not Modified Modified
Right 231 2
Wrong 4 22
5.4 RAG
Figure 13: Retrieval Augmented Generation (RAG) can improve the performance of OptiMUS-0.3
Easy Hard
OptiMUS-0.3 88.6% 73.7%
RAG on Clauses 83.8% 79.6%
RAG on Problem Description 86.7% 45.5%
RAG on Problem Labels 87.1% 53.6%
Best 88.6% 79.6%
Retrieval Augmented Generation (RAG) enhances the performance of the LLMs by augmenting the prompt with specialized data relevant to the task. It has shown success in Q&A systems that require domain-specific knowledge Lewis et al. (2020), Gao et al. (2023b), Liu et al. (2020). The RAG module in OptiMUS-0.3 aims to improve the ability of the system to correctly model objectives and constraints by retrieving examples of similar constraints and their formulations and including them in the LLM prompt. The RAG-dataset consists of the 218 problems from the dev and test sets for which OptiMUS produced a correct formulation. For a given input problem, the RAG module of OptiMUS searches for examples from among all problems in the RAG-dataset, excluding the current input problem.
We employ RAG in two phases of Algorithm 1: during the extraction of constraints and objectives (ExtractTargets, line 4), and during the formulation of constraints and objectives (FormulateClauses, line 7). In ExtractTargets, we retrieve the five problem descriptions from the dataset that have the highest cosine similarity to the OpenAI (text-embedding-3-large) embeddings of the problem descriptions. These retrieved descriptions, along with their corresponding natural language objectives and formulations, are prepended to the original prompt. In FormulateClauses, we identify the ten constraints or objectives that are most similar in terms of cosine similarity to the natural language descriptions of the current objective or constraint. We then prepend only these objectives or constraints, not the problem description, to the prompt. We also explored using similar problems in FormulateClauses, by adding three problem descriptions with natural language and formulation of objectives/constraints, by either comparing the similarity of the textual descriptions or Jaccard similarity of the problem labels from Figure 1. In Fig. 13, notice that for the easy problems, all of these RAG variants decrease performance, but for hard problems, RAG on individual clauses can improve performance substantially. This might suggest that if the problem is easy and the LLM is capable of solving it, adding RAG confuses it, but if it needs to employ a technique such as big-M, having similar examples can increase performance.
To understand how RAG affects the performance, we hand-crafted some examples. Fig. 14 (left) shows a RAG setup where we pass the constraint “No more than $20\%$ of the article prices can change” alongside a randomly chosen example of a correctly formulated constraint. The model fails to correctly formulate the intended constraint. Conversely, in Fig. 14 (right), the same constraint is provided along with an example demonstrating the big-M technique. This relevant example guides the model to correctly formulate the constraint.
Hence we see that a well-chosen RAG example can improve performance, while a poorly chosen example can degrade performance. We hypothesize that irrelevant examples can distract the LLM from focusing on the constraint at hand. We suspect that overcoming this hurdle will present a substantial challenge, and so we leave to future work further development of RAG for optimization modeling. An improved RAG might replace the default distance metric used in the RAG vector search, cosine similarity, with one tailored to the optimization context (including search that is agnostic to variable name and aware of constraint type).
Figure 14: RAG helps with the relevant examples can help OptiMUS model the constraints correctly (right), while using RAG with irrelevant examples can confuse OptiMUS (left).
6 Optimization Techniques in OptiMUS
Many of the advances in optimization solvers in the last several decades rely on the fact that solvers are faster when they can exploit particular structures within the optimization problem. One of the main tasks of an optimization expert is to identify these structures so as to choose an appropriate solver or parameter settings. However, detecting useful structures in a given optimization problem can be very challenging, since it requires both a deep understanding of the mathematical structure of the problem and the capabilities of existing solvers.
In view of this challenge, OptiMUS-v0.3 includes an optimization technique detection module, which we describe in this next section. The structure detection module is composed of two agents:
  • a structure detection agent optimization structures
  • a customized optimization coding agent
Figure 15: Structure pool in OptiMUS
6.1 Design of Structure Detector and Advanced Optimization Coding
OptiMUS maintains a pool of optimization structures which are commonly exploited in optimization software. To detect structure in an optimization problem, OptiMUS iterates through these structures and format them into a structure detection prompt. Within each prompt, the LLM is provided with the description of the structure, explained by an example illustrating how the structure should be exploited. The LLM is asked to decide whether the structure can be applied to the existing formulation. Upon identifying the appropriate structure, the formulation is adjusted to high-light the problem structure.
Currently, OptiMUS considers the following three types of structures.
Problem structure
These structures apply when MILP formulation is applied to solve problems that boil down to some more specific type. Customized solvers often exist for these problems and can dramatically improve solution speed. SAT, traveling salesman (TSP), vehicle routing (VRP) knapsack, network flow, constraint programming fall into this category.
Constraint (variable) structure
These structures are considered within an MILP model and are defined on a subset of variables or constraints. Optimization solvers can exploit these structures to improve performance Gamrath et al. (2016), and customized interface are often available. Using the interface not only reduces the complexity of (and potential for errors in) using auxiliary variables or constraints, but also informs the solver of the existence of structure that can be exploited to solve the problem faster. Typical examples of structure include Special Ordered Set (SOS) Beale and Forrest (1976), indicator variables, semi-continuous variables, and general constraints Bertsimas and Tsitsiklis (1997a).
Large-scale optimization technique and advanced optimization coding
The aforementioned structures are generally straight-forward to exploit: either changing the optimization solver or modifying a few lines of modeling code suffices. However, in large scale optimization, it is common that exploiting structure becomes highly nontrivial: in these applications, optimization solvers are often embedded into some high-level optimization framework and called as a subroutine. Typical examples including decomposition structure, column generation, and cutting plane methods. To leverage these structures, optimization coding needs to invoke advanced solver interface (for example, call-back, model attribute query and analysis), which goes far beyond modeling itself. OptiMUS implements a dedicated coding agent that adapts available modeling code to some user-specified framework.
To illustrate the performance of OptiMUS, we provide showcase examples to demonstrate the efficiency of OptiMUS in optimization structure detection.
Example: Indicator Constraint Definition An indicator constraint over a binary variable $z$ and a constraint $C$ states that $z=1\Rightarrow C$. Description A retail company wants to open new stores in a set of $K$ potential locations. A store at location $k$ must have a minimum staff of $n_{k}$ employees (to ensure adequate customer service) and a minimum inventory level of $l_{k}$ units (to avoid stockout). The maximum inventory level for store $k$ is $u_{k}$, and there is a fixed cost associated with opening each store. Objective The company aims to maximize the total expected revenue from the opened stores. Relevant Constraints If a store at location $k$ is open, it must have at least $n_{k}$ employees. Moreover, let the inventory level of store $k$ be $y_{k}$. Formulations Original: $x_{k}\in\{0,1\},~{}y_{k}\leq u_{k}x_{k},$ $z_{k}\geq l_{k}-M(1-x_{k})$ OptiMUS: $x_{k}$$\Rightarrow$ $y_{k}\leq u_{k}$ $x_{k}$$\Rightarrow$ $z_{k}\geq l_{k}$ Performance plot [Uncaptioned image]
Example: Special Ordered Set Definition Given a set of variables $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$, $\{x_{1},\ldots,x_{n}\}\in\operatorname{SOS}_{1}$ if at most one element of $\mathcal{X}$ can be nonzero $\{x_{1},\ldots,x_{n}\}\in\operatorname{SOS}_{2}$ if at most two elements of $\mathcal{X}$ can be nonzero Description An airline needs to assign $N$ crew members to a set of $K$ flights for a day. Each crew member can only be assigned to at most one flight at a time, and there must be a minimum rest period of 2 hours between consecutive flights to comply with aviation regulations and ensure crew well-being. Each crew member has a maximum number of flying hours per day, set at 8 hours, to comply with aviation regulations and prevent fatigue. There are minimum and maximum layover times between flights: minimum layover of 2 hours (to allow for crew rest and flight preparations) and maximum layover of 6 hours (to optimize crew utilization and reduce idle time). Assigning a crew member to a specific flight has some cost. Objective Minimize the total crew cost while ensuring that each flight has the required number of qualified crew members, satisfying the rest period constraints, maximum flying hours per crew member, and layover time constraints. Relevant Constraints Each crew member can work on only one flight. Let $x_{ij}\in\{0,1\}$ denote whether crew $i$ works on flight $j$ and $T_{ij}$ denote the working time of crew $i$ on flight $j$. Formulations Original: $\textstyle\sum_{j=1}^{K}x_{ij}\leq 1,\quad$ $\forall i\in[N],$ $T_{ij}\leq 24x_{ij},\quad$ $\forall i\in[N],j\in[K]$ OptiMUS: $\{T_{i1},\ldots,T_{iK}\}\in\operatorname{SOS}_{1}$ Performance plot [Uncaptioned image]
6.1.1 Large-scale column/constraint generation
Consider a general optimization problem
$\begin{array}[]{ll}\mbox{minimize}&c(x)\\ \text{subject to}&x\in\mathcal{X}_{k},~{}k=1,\ldots,m\end{array}$ ($\mathcal{P}$)
with variable $x\in\mathbb{R}^{n}$ in which either $m$ or $n$ is large. Column and constraint generation methods, respectively, allow for scalable solutions to such problems. For example, constraint generation for $m\gg n$ proceeds by first restricting optimization problem to a subset $\mathcal{W}\subseteq[m]$ of constraints, often chosen at random: define $x_{\mathcal{W}}^{\star}$ to be the solution of the problem
$\begin{array}[]{ll}\mbox{minimize}&c(x)\\ \mbox{subject to}&x\in\mathcal{X}_{k},~{}k\in\mathcal{W}\end{array}$ ($\mathcal{P}_{\mathcal{W}}$)
with variable $x\in\mathbb{R}^{n}$. When the number of constraints $|\mathcal{W}|$ is small, the resulting problem ($\mathcal{P}_{\mathcal{W}}$) can be solved more efficiently. Given the solution $x_{\mathcal{W}}^{\star}$ to ($\mathcal{P}_{\mathcal{W}}$), constraint generation solves a pricing problem to find the most violated constraint
[…]
This constraint $j_{\mathcal{W}}$ is added to the set $\mathcal{W}$ and the algorithm continues until convergence. As the algorithm continues until all constraints are satisfied, the solution to the final restricted master problem ($\mathcal{P}_{\mathcal{W}}$) is also optimal for the original problem ($\mathcal{P}$).
We provide an example of constraint generation in OptiMUS in Fig. 16 on a security-constrained unit commitment problem. For details of the problem see the appendix.
Figure 16: Left: constraint generation prompt. Right: performance plot of solving to 5% gap.
7 Conclusion
How can we leverage LLMs to achieve complex goals? This paper interrogates this question in the domain of optimization and showcases the importance of modular structure. We develop OptiMUS-0.3, a modular LLM-based agent designed to formulate and solve optimization problems from natural language descriptions. Our research serves as a proof-of-concept, illustrating the potential for automating various stages of the optimization process by combining LLMs with traditional solvers. To showcase the performance of OptiMUS-0.3, we released NLP4LP, a comprehensive dataset of long and challenging optimization problems to demonstrate the efficacy of the techniques implemented within OptiMUS-0.3. OptiMUS-0.3 achieves SOTA performance across all existing datasets, and scales to problems with large amounts of data and long descriptions.
Real-world optimization problems are often complex and multifaceted. Developing LLM-based solutions for these problems requires domain-specific considerations, including integrating existing optimization techniques to leverage problem structure. We are at the early stages of this research, but anticipate significant developments that will enable these systems to address more complex, industrial-level problems. It is interesting to notice that the challenge of using AI for an applied domain is much larger in safety-critical domains such as self-driving, which demand extremely high accuracy, than in domains where AI can function as an assistant and where answers are easy to check, as in theorem-proving or optimization. Here, AI systems with moderate accuracy can still usefully augment human effort.
Future directions.
The development of OptiMUS opens up several promising avenues for future research in AI-assisted optimization modeling. Enhancing LLMs’ capacity to interpret nuanced or ambiguous problem descriptions could improve its robustness and applicability, potentially through specialized language models fine-tuned on optimization terminology. An interesting area of exploration could be the combination of structured pipelines and more flexible agent-based systems, shedding light on their relative strengths and integration potential. The development of mechanisms for efficiently integrating user feedback and iteratively refining models presents another intriguing research direction. Exploring the applicability of OptiMUS’s modular LLM structure to other complex problem-solving tasks beyond optimization might yield insights into building more capable AI systems for a range of applications. Finally, the creation of more comprehensive benchmarks for assessing AI-assisted optimization modeling tools could drive progress in this field.
Appendix: SCUC problem
Description.
The Security Constrained Unit Commitment (SCUC) is a critical optimization problem in the operation of electrical power systems. Its goal is to schedule generation units in a cost-effective manner while ensuring reliable operation over a specified time period (typically 24 hours), considering the system’s physical and operational constraints. This appendix presents a detailed description of the SCUC problem, including its objectives and constraints:
Objective.
The objective of the SCUC problem is to minimize the total operational cost of the power system. This cost generally includes:
  • Fuel Cost: The cost associated with consuming fuel to generate electricity.
  • Start-up and Shut-down Cost: Cost incurred from starting up or shutting down generation units.
  • Emission Cost: Cost related to the emissions produced by the generating units, if applicable.
Constraints.
The SCUC must satisfy several constraints to ensure the safe and reliable operation of the power system:
  • Power Balance Constraint: Ensures that the total power generation meets the total demand plus system losses at every interval. Mathematically, this is expressed as the sum of the outputs of all online generators being equal to the sum of the demand and the transmission losses.
  • Minimum and Maximum Output Limits: Each generator has a minimum and maximum generation capacity when it is online. The SCUC ensures that the output of each unit stays within these bounds.
  • Ramp-up and Ramp-down Limits: These limits specify the maximum rate at which a generator can increase or decrease its output. They are critical for handling load changes throughout the day.
  • Unit Start-up and Shut-down Constraints: These constraints handle the logistics of turning units on or off. Start-up constraints may include minimum down times (the minimum time a unit must remain off before it can be restarted) and minimum up times (the minimum time a unit must remain on once started).
  • Minimum Up and Down Time Constraints: Ensures that once a generator is turned on, it stays on for at least its minimum up time, and similarly, once turned off, it remains off for at least its minimum down time. These constraints are crucial for the mechanical integrity of the generation units.
  • Reserve Requirements: The system operator must ensure that sufficient spinning and non-spinning reserves are available. These reserves are needed to handle sudden increases in demand or unexpected generator failures.
Appendix: Checking formulation equivalence
To evaluate the correctness of the formulation of a given optimization problem, OptiMUS evaluates the formulation on a numerical example and checks that the optimal value matches the ground truth and that all constraints are satisfied. This approach runs the risk that an incorrect formulation might return the correct solution on the particular problem data chosen, but not for different problem data. An alternative approach would be to compare the formulation against a reference ground-truth mathematical formulation, and evaluate whether the formulations are the same. This alternative approach has the benefit that it is independent of numerical data, and indeed would not require the availability of any problem data. This benefit is important for complex and realistic industrial problems, for which true problem data may be proprietary and feasible problem data may be difficult to generate.
Unfortunately, there are several serious obstacles to comparing the formulation of a problem against a reference formulation. First, for most reasonable notions of equivalence, two formulations can both be correct even if they are not equivalent. Worse, the problem is NP-hard: in its general form, the problem of checking the equivalence of formulations can be reduced to (a general form of) the well-known Graph Isomorphism Problem (GIP). In the discussion that follows, we restrict our attention to feasibility problems for simplicity; the extension to more general optimization problems is similar.
Formulation equivalence
Two mathematical formulations are considered equivalent if there exists a one-to-one correspondence between their variables and constraints in which the coefficients of corresponding pairs of variables are equal in every corresponding pair of constraints. In other words, the coefficients of each variable, whether symbolic or numeric, must align under the mapping. Note that it is possible for two formulations to describe the same feasible set without such a correspondence, in which case they are not considered equivalent under this strict definition.
Reducing formulation equivalence to generalized graph isomorphism
To compare two formulations, we first represent each formulation as a bipartite graph $G$ where variables and constraints form separate sets of nodes. A variable node is linked to a constraint node by an edge if the variable appears in the constraint, with the edge labeled by the variable’s coefficient in the constraint. This unique graph representation allows us to show the equivalence of two formulations by establishing an isomorphism between their respective graphs.
Given two formulations, $\mathcal{F}$ and $\mathcal{F}^{\prime}$, with differing variables and constraints, we construct their corresponding graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$. The challenge then is to find a correspondence between the graphs’ nodes that respects the connectivity and coefficient labels, effectively determining graph isomorphism between $\mathcal{G}$ and $\mathcal{G}^{\prime}$.
Reducing graph isomorphism to formulation equivalence
Consider two graphs $\mathcal{G}_{1}=(V_{1},E_{1})$ and $\mathcal{G}_{2}=(V_{2},E_{2})$. To check for isomorphism, we construct corresponding mathematical formulations, $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.
Constructing mathematical formulations
  1. 1.
    Variables and constraints: For each vertex $v$ in $V_{i}$, define a variable $x_{v}$. For each edge $(u,v)\in E_{i}$, create a constraint $x_{u}+x_{v}=1$.
  2. 2.
    Formulation: Each formulation $\mathcal{F}_{i}$ includes:
    • Variables: $\{x_{v}\mid v\in V_{i}\}$
    • Constraints: $\{x_{u}+x_{v}=1\mid(u,v)\in E_{i}\}$
An algorithm that finds a bijective mapping between the variables and constraints of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ effectively establishes a one-to-one correspondence between the vertices and edges of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
This mapping ensures:
  • Each variable in $\mathcal{F}_{1}$ maps to a unique variable in $\mathcal{F}_{2}$, corresponding to a vertex in $\mathcal{G}_{2}$.
  • Each constraint in $\mathcal{F}_{1}$ maps to a unique constraint in $\mathcal{F}_{2}$, corresponding to an edge in $\mathcal{G}_{2}$.
Thus, verifying the equivalence of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ is equivalent to determining if $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are isomorphic, showing that this problem is as computationally complex as the Graph Isomorphism Problem.
Concluding the reduction
This approach effectively reduces the problem of checking graph isomorphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ to verifying the equivalence of the mathematical formulations $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. This reduction not only underscores the computational complexity of formulation equivalence but also classifies it within the same complexity category as the (NP-complete) Graph Isomorphism Problem.